# LLMs’ Reshaping of People, Processes, Products, and Society in Software Development: A Comprehensive Exploration with Early Adopters

### v1

@@ 3.3.2. Optimal Use Cases
  • Better Text than Code Generations: Two participants observed that LLMs seemed to perform better at generating textual content than code, possibly due to their training data composition (Wu et al., 2023). P8 shared that, in his experience, “[ChatGPT] works better for text, not the codebase.
3.3.3. Security
  • Providing Sufficient Security: Thirteen participants expressed that they considered LLMs’ code to be sufficiently secure for their purposes, primarily because they were not using it in production environments or for critical applications. It was also mentioned that LLMs’ code is as secure as any code publicly available. Seven participants shared that this was because of their specific use case, like P2, who noted that “I’ve never had to use it for any security aspects.” P14 additionally used a VPN, and P4 and P15 shared that they believed their reading of the generated code was another layer of added protection against security concerns (P4: “It’s not like I’m not reading the code. So if it’s trying to like, wipe my hard drive, or leak customer data — There’s no scenario where that could happen.”)
@@ 3.3.3. Security
  • Concerns Sending Data to LLMs: Eight participants shared that they specifically did not provide LLMs with identifying, confidential, or otherwise proprietary information in their prompts. P3 expressed concerns over the lack of “trust and safety controls.” A few participants noted concerns about their data being used by companies like OpenAI. P4 stated, “There is, of course, concern with what code am I sending to OpenAI.” P12 shared related concerns: “If those interactions with Bard or ChatGPT get logged and used for training data in the future, it could be that those models start outputting production code or like, our own internal code. And that’s something we want to avoid.” It is worth noting that, since April 2023, ChatGPT has included an option to prevent a user’s queries from being used to train or improve the model (Dastin and Tong, 2023)—the feature that both P14 and P15 mentioned that they used in order to reduce their security concerns. At present, Google Gemini additionally allows users some control over how their data is used by Google (n.a., 2024c), and Copilot allows users to opt out of sharing their data (n.a., 2024b).
@@
  • Concerns using Data from LLMs: Eight participants emphasized developers’ responsibility to ensure code correctness, especially regarding the possibility copying and using code without review. Two participants explicitly stated that LLM code should not be copied without review due to security concerns; P11 shared that, “I think it’s dangerous, right? Like, if out of our work on a team, somebody would just copy-paste ChatGPT [code], you know, I’ll probably be annoyed by it.” P1 further expressed concerns over malicious actors potentially injecting poisoning code that goes on to then further train ChatGPT, which could result in the tool generating exploitable code in the future.
    We note that both this and the above developer-identified challenge are clearly informed by applying knowledge of how LLMs work to determine possible risks of sharing data with LLMs and using LLM outputs in production code.
3.4. RQ4: How may the software industry and education be affected by LLMs?
@@
The key themes that emerged in understanding how LLMs impact two areas of society—the software industry and CS education—are as follows:
3.4.1. Industry
Figure 5. Findings related to RQ4, split horizontally by findings on industry (above) and CS education (below) and vertically by opportunities (left) and challenges (right).
@@ 3.4.1. Industry
  • Comparing LLMs to Existing Entities or Roles: We found nine interviews in which LLMs were fulfilling roles commonly filled by other people or tools, including as a pair-programmer, assistant or secretary, junior developer, rubber duck, or simply a tool. For instance, P10, whose colleagues had been all laid off, mentioned, “I’ve been working solo by myself for a few months now. [ChatGPT] is my junior who [is] trying to help me.” He further added, “I still write [the code] the way I do. [ChatGPT] just has some extra eyes that [are] helping me, to guide me through [the coding process] or to give me some recommendations overall.” P5 also mentioned, “[ChatGPT] is like the rubber ducky that we would have, except now it produces answers, and it talks to me, and it gives me all of the advice that I would need.” These results suggest that developers are in the process of figuring out the future of their careers and how LLMs may ultimately fit in.
@@
  • Minimal Impact on the Job Market: Nine participants expressed that jobs in the software development field would remain largely unaffected. They argued that software development involves more than just code writing, and LLMs lack the capability to fully substitute human developers. As P3 noted, “I don’t really see it replacing people because I think if you’re at the point, at least with the current version of it, that it could do your job, you probably weren’t as good of a programmer in the first place. .”
@@
  • Widespread Use Among Peers: Four participants observed widespread use of ChatGPT among their peers. P7, for example, pointed out, “[…] maybe 90%, maybe 80% of my team started using ChatGPT from January or February [2023]. Maybe because we’re a tech company, people know this quicker than the others. And now I think 100% of people are starting to use it, just depending on the use case.” This suggests a growing acceptance and adoption of LLMs within professional environments, particularly within tech companies.
@@
  • Lowering Entry Barriers: Three participants mentioned LLMs could lower the barrier for entry-level positions by providing assistance and resources. P6 shared his hopes for LLMs being able to answer novice programmers’ questions: “[ChatGPT] has lowered the barrier to entry in terms of coding. So you can really go and ask and stuff, and depending on what your level of expertise is or what your level of questioning is, it will provide you with the level of answers. So you don’t have to rely on books or universities and then going through the [usual] courses. Not that there is anything wrong with them, but it’s a different way of approaching programming.
@@
  • Absence and Necessity of Guidelines: Eight participants emphasized the importance of establishing guidelines for LLM use, especially for larger companies or those dealing with security-sensitive information. They highlighted the need for clear rules to ensure the secure usage of LLMs. P9, for example, noted, “I think 100% [that companies should have guidelines], especially when you’re dealing with things that are sensitive; for example, financial institution or medical [data]. And I actually do think that instead of them trying to use third-party LLMs from, for example, OpenAI, they should start building their own models and have them housed within their own PVC cloud so that it’s secure enough for their standard instead of trying to buy this off-shelf.” P7 highlighted the utility of more specific guidelines, noting that some companies should provide prompt templates. As he noted, “I think that would be good [to have guidelines] because I’ve heard there are some question templates to let ChatGPT give us what we want.
    Of these eight participants, five noted that their companies lacked specific guidelines for LLM use. However, they noted that existing rules, such as not sharing sensitive information, apply to using LLMs. P5 shared, “We’re a pretty small company, so we don’t have too [many] regulations against it, and a lot of people work really like me and the guy that I’m working under. He uses it fairly heavily as well. So, outside of making sure that we don’t provide any sensitive data or information to ChatGPT, we’re pretty much just okay to use it.” Similarly, P2 stated, “We had these guidelines even before ChatGPT; the guidelines were that we are not allowed to share any client-specific information with anyone.
  • Lowering the Need for Certain Roles: Five participants highlighted LLMs’ potential to decrease the number of developers required for certain tasks and roles, such as user interface professionals or data scientists. P8 mentioned, “I’m not sure [if] it’s changed a lot of things in software engineering, but in the other fields, like data science, I’m seeing a [change] because, in those fields, there are people who don’t have a lot of deep knowledge about computer science. […] For example, they have a piece of Python code, and they want to do some logic. They have a matrix, and they want to transpose it. [So] they can use ChatGPT very easily, and it does that for them.
  • Undermining Entry-Level Positions: LLMs’ potential to undermine the demand for entry-level roles emerged three times in our study. P6, for instance, mentioned, “I do find the level of code generated is sometimes almost as good as a junior software developer. So I think that really up the bar of hiring for junior software developers. When I graduated, ChatGPT wasn’t around, and I probably didn’t have to expect to know that much.” P15 also stated that newer developers should broaden their skill set in order to stay ahead of technological advancements and mitigate their negative impacts:“As long as you have a bigger picture of things, and you’re able to engineer things more thoroughly, you can beat the pace of ChatGPT.
  • Exploiting LLMs in Job Interviews: Two participants raised concerns about candidates using LLMs to pass interviews without genuine knowledge or skills, so formulaic interview techniques might need to change. P1, for example, mentioned that people could very easily use ChatGPT to pass, “ without actually knowing how to pass the interview. I don’t think that necessarily makes them a bad software engineer, but I think it can break this formulaic interview a little bit.
3.4.2. Education
@@
  • Impracticality of Bans: Three participants viewed the issue from another perspective: the impracticality of banning it. For instance, P2 stated, “You cannot run away from it because everyone has access […] Students are going to use it, no matter [what].” Participants emphasized the potential benefits of LLMs in supplementing learning. As P16 shared, “I think degree programs should encourage students to use ChatGPT in the right way.” Participants felt this could be particularly important for underprivileged students who may lack access to traditional tutoring or resources.
@@
  • Emphasizing Fundamental Concepts: Six participants stressed that foundational knowledge, theoretical understanding, and fundamental concepts should be prioritized over the integration of ChatGPT. They believed that topics like software architecture, algorithm design, and problem-solving skills should take precedence. For instance, P1 mentioned, “I think [universities] should be doing more of what they should have been doing in the first place, which is not necessarily focusing on specific implementations and specific kind of optimal algorithms, but rather on the bigger, more difficult-to-grasp ideas that have to do with architecting software and that have to do with thinking about what it takes to get from a business idea to an actual product.
    Concerns over LLMs being just the latest tool or technology that may become obsolete also appeared. For example, P6 mentioned, “One thing remains constant [in CS education] that you are studying and you’re understanding and you’re learning the programming languages and you’re learning the technology that will be obsolete by the time you get to the market for a job. I don’t think that the programs should [change based] on what’s the latest and the greatest, ChatGPT, Anthropic AI, this AI, that AI, or this fancy new programming language. I think that’s the wrong way to go. I think, in general, the program should be focused on more fundamentals because those largely remain the same.

### v2

@@ 4.3.2 Optimal Use Cases
  • Better at Text than Code: Two participants observed that LLMs seemed to perform better at generating textual content than code, possibly due to their training data composition [67].
4.3.3 Security
  • Providing Sufficient Security: Thirteen participants expressed that they considered LLMs’ code to be sufficiently secure for their purposes, primarily because they were not using it in production environments or for critical applications. It was also mentioned that LLMs’ code is as secure as any publicly available code. Seven participants shared that this was because of their specific use case, like P2, who noted that “I’ve never had to use it for any security aspects.” P14 additionally used a VPN, and P4 and P15 shared that they believed their reading of the generated code was another layer of added protection against security concerns.
@@ 4.3.3 Security
  • Concerns Sending Data to LLMs: Eight participants shared that they specifically did not provide LLMs with identifying, confidential, or otherwise proprietary information in their prompts. P3 expressed concerns over the lack of trust and safety controls.” A few participants noted concerns about their data being used by companies like OpenAI. P12 shared related concerns: If those interactions with Bard or ChatGPT get logged and used for training data in the future, it could be that those models start outputting production code or like, our own internal code. And that’s something we want to avoid.” It is worth noting that, since April 2023, ChatGPT has included an option to prevent a user’s queries from being used to train or improve the model [68]; the feature that both P14 and P15 mentioned that they used in order to reduce their security concerns. At present, Google Gemini additionally allows users some control over how their data is used by Google [69], and Copilot allows users to opt out of sharing their data [70].
@@
  • Concerns using Data from LLMs: Eight participants emphasized developers’ responsibility to ensure code correctness, especially regarding the possibility of copying and using code without review. Two participants explicitly stated that LLM code should not be copied without review due to security concerns; P11 shared that, “I think it’s dangerous, right? Like, if out of our work on a team, somebody would just copy-paste ChatGPT [code], you know, I’ll probably be annoyed by it.” P1 further expressed concerns over malicious actors potentially injecting poisoning code that goes on to then further train ChatGPT, which could result in the tool generating exploitable code in the future. We note that both this and the above developer-identified challenge are clearly informed by applying knowledge of how LLMs work to determine possible risks of sharing data with LLMs and using LLM outputs in production code.
4.4 RQ4: How may the software industry and education be affected by LLMs?
@@
The key themes that emerged in understanding how LLMs impact two areas of society (the software industry and CS education) are as follows:
4.4.1 Industry
@@
  • Comparing LLMs to Existing Entities or Roles: We found nine interviews in which LLMs were fulfilling roles commonly filled by other people or tools, including as a pair-programmer, assistant or secretary, junior developer, rubber duck, or simply a tool. For instance, P10, whose colleagues had been all laid off, mentioned, I’ve been working solo by myself for a few months now. [ChatGPT] is my junior who [is] trying to help me.” P5 also mentioned, [ChatGPT] is like the rubber ducky that we would have, except now it produces answers, and it talks to me, and it gives me all of the advice that I would need.” These results suggest that developers are in the process of figuring out the future of their careers and how LLMs may ultimately fit in.
@@
  • Minimal Impact on the Job Market: Nine participants expressed that jobs in the software development field would remain largely unaffected. They argued that software development involves more than just code writing, and LLMs lack the capability to fully substitute human developers. As P3 noted, “I don’t really see it replacing people because I think if you’re at the point, at least with the current version of it, that it could do your job, you probably weren’t as good of a programmer in the first place.
@@
  • Widespread Use Among Peers: Four participants observed widespread use of ChatGPT among their peers, suggesting a growing acceptance and adoption of LLMs within professional environments, particularly within tech companies.
@@
  • Lowering Entry Barriers: Three participants mentioned LLMs could lower the barrier for entry-level positions by providing assistance and resources. P6 shared his hopes on LLMs being able to answer novice programmers’ questions: “[ChatGPT] has lowered the barrier to entry in terms of coding. So you can really go and ask and stuff, and depending on what your level of expertise is or what your level of questioning is, it will provide you with the level of answers.
@@
  • Absence and Necessity of Guidelines: Eight participants emphasized the importance of establishing guidelines for LLM use, especially for larger companies or those dealing with security-sensitive information. P9, for example, noted, “I think 100% [that companies should have guidelines], especially when you’re dealing with things that are sensitive; for example, financial institution or medical [data]. And I actually do think that instead of them trying to use third-party LLMs from, for example, OpenAI, they should start building their own models and have them housed within their own PVC cloud so that it’s secure enough for their standard.” Of these eight participants, five noted that their companies lacked specific guidelines for LLM use, though existing rules about not sharing sensitive information applied.
  • Lowering the Need for Certain Roles: Five participants highlighted LLMs’ potential to decrease the number of developers required for certain tasks and roles, such as user interface professionals or data scientists.
  • Undermining Entry-Level Positions: LLMs’ potential to undermine the demand for entry-level roles emerged three times in our study. P6, for instance, mentioned, “I do find the level of code generated is sometimes almost as good as a junior software developer. So I think that really up the bar of hiring for junior software developers.” P15 also stated that newer developers should broaden their skill set in order to stay ahead of technological advancements.
  • Exploiting LLMs in Job Interviews: Two participants raised concerns about candidates using LLMs to pass interviews without genuine knowledge or skills, suggesting that formulaic interview techniques might need to change.
4.4.2 Education
@@
  • Impracticality of Bans: Three participants viewed the issue from another perspective: the impracticality of banning it. For instance, P2 stated, “You cannot run away from it because everyone has access […] Students are going to use it, no matter [what].” Participants emphasized the potential benefits of LLMs in supplementing learning, particularly for underprivileged students who may lack access to traditional tutoring or resources.
@@
  • Emphasizing Fundamental Concepts: Six participants stressed that foundational knowledge, theoretical understanding, and fundamental concepts should be prioritized over the integration of ChatGPT. They believed that topics like software architecture, algorithm design, and problem-solving skills should take precedence. For instance, P1 mentioned, “I think [universities] should be doing more of what they should have been doing in the first place, which is not necessarily focusing on specific implementations and specific kind of optimal algorithms, but rather on the bigger, more difficult-to-grasp ideas that have to do with architecting software and that have to do with thinking about what it takes to get from a business idea to an actual product.” Concerns over LLMs being just the latest tool or technology that may become obsolete also appeared. For example, P6 mentioned, “One thing remains constant [in CS education] that you are studying and you’re understanding and you’re learning the programming languages and you’re learning the technology that will be obsolete by the time you get to the market for a job. […] I think, in general, the program should be focused on more fundamentals because those largely remain the same.